# second generation 3G third generation

### v1

@@ VI Conclusion
    In this paper, we proposed a novel tensor-based semi-blind receiver design for BD-RIS-assisted multiple-input multiple-output (MIMO) communication systems, where we assume a time-varying UT- intelligent reconfigurable surface (IRS) channel, which leads us to a fourth-order PARATUCK tensor model for the received signals. Exploiting this, we formulated a semi-blind receiver that is a data-aided channel estimator, avoiding the use of pilot sequences, while jointly estimating the BD-RIS-BS channel, the UT-BD-RIS channel, and the transmitted symbols. Specifically, we proposed to the recipients: I) The PAKRON receiver (Algorithm 1) is a two-stage approach where, first, the data symbols are estimated jointly with a structured matrix $\boldsymbol{\Omega}$ which is comoposed by the Kronecker product between the BD-RIS-BS and the data matrix symbols in an iterative way by means of a BALS. Next, in the second stage, a Kronecker factorization is applied on the estimated version of the $\boldsymbol{\Omega}$ to estimate the BD-RIS-BS channel component as well as the symbols. II) The TUCKER receiver is a fourth-order tensor decomposition method, where the channel and the data are jointly and iteratively estimated in a single stage via an TALS algorithm (4).
    Our results indicate that the TUCKER receiver is more efficient than the PAKRON receiver in terms os channel estimation and data symbol accuracy. However, it is more computationally complex than PAKRON. Additionally, both of our methods outperform the competing method LS, as it does not exploit the inherent multidimensional structure of the received signal. On the other hand, the competing method LS-KRF [^7] The same for BTKF and BTALS in [Sokal_BD_RIS_2024]. exploit the tensor structure to surpass our proposed method. However, the competing method is a pilot-assisted scheme, whereas our approach is a semi-blind scheme suitable for scenarios with limited or unavailable pilot resources. A possible hybrid implementation of the proposed semi-blind receiver uses the soft estimates of the data symbol matrix delivered by PAKRON as an initialization for the TUCKER receiver. Another perspective of this work is to investigate the optimal design of the matrix $\boldsymbol{\Psi}$. Additionally, new tensor structures as well as the Cramér-Rao bound can be derived for the different receivers.

### v2

@@ VI Conclusion
    In this paper, we proposed a novel tensor-based semi-blind receiver design for BD-RIS-assisted multiple-input multiple-output (MIMO) communication systems, where we assume a time-varying UT- intelligent reconfigurable surface (IRS) channel, which leads us to a fourth-order PARATUCK tensor model for the received signals. Exploiting this, we formulated a semi-blind receiver that is a data-aided channel estimator, avoiding the use of pilot sequences, while jointly estimating the BD-RIS-BS channel, the UT-BD-RIS channel, and the transmitted symbols. Specifically, we proposed to the recipients: I) The PAKRON receiver (Algorithm 1) is a two-stage approach where, first, the data symbols are estimated jointly with a structured matrix $\bm{\Omega}$ which is comoposed by the Kronecker product between the BD-RIS-BS and the data matrix symbols in an iterative way by means of a BALS. Next, in the second stage, a Kronecker factorization is applied on the estimated version of the $\bm{\Omega}$ to estimate the BD-RIS-BS channel component as well as the symbols. II) The TUCKER receiver is a fourth-order tensor decomposition method, where the channel and the data are jointly and iteratively estimated in a single stage via an TALS algorithm (4).
    Our results indicate that the TUCKER receiver is more efficient than the PAKRON receiver in terms os channel estimation and data symbol accuracy. However, it is more computationally complex than PAKRON. Additionally, both of our methods outperform the competing method LS, as it does not exploit the inherent multidimensional structure of the received signal. On the other hand, the competing method LS-KRF [^7] The same for BTKF and BTALS in [Sokal_BD_RIS_2024]. exploit the tensor structure to surpass our proposed method. However, the competing method is a pilot-assisted scheme, whereas our approach is a semi-blind scheme suitable for scenarios with limited or unavailable pilot resources. A possible hybrid implementation of the proposed semi-blind receiver uses the soft estimates of the data symbol matrix delivered by PAKRON as an initialization for the TUCKER receiver. Another perspective of this work is to investigate the optimal design of the matrix $\bm{\Psi}$. Additionally, new tensor structures as well as the Cramér-Rao bound can be derived for the different receivers.